if 2 hotdogs and 1 hamburger costs 3.39. and 3 hotdogs and 3 hamburgers cost 6.57, how much does a hotdog cost?
Answer by checkley71(8403)

2D+1B=3.39
3D+3B=6.57 MULTIPLY EQ #1 BY -3 WE GET
-6D-3B=-10.17 NOW ADD TO EQ #2
3D+3B=6.57
-------------------------
-3D=360
D=3.60/-3
D=$1.20 THE COST OF ONE HOTDOG